Bill and Ted are walking on a trail. For the first 10.0 minutes of the walk, they are full of energy and travel at an average speed of 2.23 meters per second. They then slow their pace to 1.75 meters per second for the remaining 50 minutes of their walk. What is their average velocity?


1
Expert's answer
2020-12-21T10:39:34-0500
"v=\frac{v_1t_1+v_2t_2}{t_1+t_2}\\v=\frac{10(2.23)+50(1.75 )}{10+50}=1.83\frac{m}{s}"